I have been able to determine a. but am stuck on b. and c. for this question...

If Wild Widgets, Inc., were an all-equity company, it would have a beta of 0.9. The company has a target debt–equity ratio of .4. The expected return on the market portfolio is 12 percent, and Treasury bills currently yield 4.1 percent. The company has one bond issue outstanding that matures in 20 years and has a coupon rate of 7.2 percent. The bond currently sells for $1,090. The corporate tax rate is 35 percent.

a. What is the company’s cost of debt? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of debt__ 6.39% __

b. What is the company’s cost of equity? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) Cost of equity ____% _

c. What is the company’s weighted average cost of capital? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) WACC____ %_
Answer #1

Answer b.

Unlevered Beta = 0.90

Levered Beta = Unlevered Beta * [1 + (Debt-Equity Ratio) * (1 - Tax Rate)]
Levered Beta = 0.90 * [1 + 0.40 * (1 - 0.35)]
Levered Beta = 1.134

Cost of Equity = Risk-free Rate + Levered Beta * (Market Return - Risk-free Rate)
Cost of Equity = 4.10% + 1.134 * (12.00% - 4.10%)
Cost of Equity = 4.10% + 1.134 * 7.90%
Cost of Equity = 13.06%

Answer c.

Weight of Debt = 0.40 / 1.40
Weight of Equity = 1.00 / 1.40

WACC = Weight of Debt * Cost of Debt * (1 - Tax Rate) + Weight of Equity * Cost of Equity
WACC = (0.40 / 1.40) * 6.39% * (1 - 0.35) + (1.00 / 1.40) * 13.06%
WACC = 1.19% + 9.33%
WACC = 10.52%